Convergence test for "\displaystyle\sum_{n=1}^\infty \frac{sin(n)}{n}".


1
Expert's answer
2021-11-12T01:23:59-0500

Let us show that the series "\sum\limits_{n=1}^{\infty}\frac{\sin(n)}n" is divergent.

Note that "|\sin x|⩾\frac{1}2" for "x∈I_k=[\frac{π}6+kπ,π−\frac{π}6+kπ]."

Let us find the length of every "I_k:"

"|I_k|=π−\frac{2π}6=\frac{2π}3>2."

We conclude that every "I_k" contains at least one natural number "n_k."

Then

"\sum\limits_{n=1}^N\frac{|\sin(n)|}n
⩾\frac{1}2\sum\limits_{n_k\le N}\frac{1}{n_k}\to\infty" when "N\u2192\u221e"

since the harmonic series "\sum\limits_{n=1}^{∞}\frac{1}n" diverges.